STEPS TOWARD EXPERIMENTAL DETECTION OF MOLECULAR PARITY VIOLATION: ROVIBRA-TIONAL ANALYSIS OF THE CF-STRETCHING MODE AND FIRST OVERTONE OF CHFBrI

\underline{S. ALBERT}, K.K. ALBERT, M. QUACK, PHYSICAL CHEMISTRY, ETH ZÜRICH, CH-8093 ZÜRICH, SWITZERLAND; S. BAUERECKER, INSTITUT FÜR TECHNISCHE UND PHYSIKALISCHE CHEMIE, TU BRAUNSCHWEIG, D-38106 BRAUNSCHWEIG, GERMANY.

The experimental detection of molecular parity violation^{a} Δ_{P V} E is of great interest because of its importance in the understanding of fundamental aspects of molecular dynamics and symmetries. One possible method for this is measuring the rovibrational or rotational frequency shifts in the infrared or microwave spectra of enantiomers^{b}. A value of 50 mHz is predicted for vibrational frequency shifts in CHFBrI^{c}, which is almost within the range of current infrared^{d} and submm wave spectroscopic resolution. We report here IR-spectroscopic results and a first high resolution analysis of the infrared spectrum of CHFBrI. The FTIR spectrum of CHFBrI was recorded at 190 K and 295 K in the regions 600–1300 cm^{−1} and 1800–2350 cm^{−1}^{e}. We were able to analyse the rovibrational spectra of CHF^{79}BrI (ν_{0} = 1060.81569 cm^{−1}) and CHF^{81}BrI (ν_{0} = 1060.77864 cm^{−1}) in the CF-stretching (ν_{6}) and its overtone regions with band centers ν_{0} = 2103.75730 cm^{−1} for CHF^{79}BrI and ν_{0} = 2103.65974 cm^{−1} for CHF^{81}BrI. We will discuss a possible application of CO_{2} laser quasi-resonant two photon transitions in the overtone region of CHFBrI. Finally, we will show submm spectra of CHFBrI recorded with the Zürich-FASSST spectrometer and discuss how submm wave spectroscopy based on FASSST^{f} and phase-locked backward wave oscillators^{g} can be used to determine line shifts on the order of mHz.

aM. Quack, J. Stohner and M. Willeke, Annu. Rev. Phys. Chem. 2008, 59, 741, A. Bakasov, T.K. Ha and M. Quack, J. Chem. Phys. 1998, 109, 7263, R. Berger and M. Quack, J. Chem. Phys. 2000, 112, 3148.

bM. Quack and J. Stohner, PRL 2000, 84, 3807, M. Quack and J. Stohner. J. Chem. Phys. 2003, 119, 11228.

cJ.K. Laerdahl, P. Schwerdtfeger and H.M. Quiney, PRL 2000, 84, 3811, R. Berger and J.L. Stuber, Mol. Phys. 2007, 105, 41.

dC. Daussy, T. Marrel, A. Amy-Klein, C. Nguyen, C. Borde and C. Chardonnet, Phys. Rev. Lett. 1999, 83, 1554

eS. Albert and M. Quack, ChemPhysChem 2007, 8, 1271.

fD. T. Petkie, T. M. Goyette, R. P. A. Bettens, S. P. Belov, S. Albert, P. Helminger and F. C. De Lucia, Rev. Sci. Instr. 1997, 68, 1675.

gF. Lewen, R. Gendriesch, I. Pak, D.G. Paveliev, M. Hepp, R. Schieder and G. Winnewisser, Rev. Sci. Instrum. 1998, 69, 32.